# The nature of ASASSN-24fw’s occultation: modelling the event as dimming by optically thick rings around a sub-stellar companion

### v1

@@ Abstract
ASASSN-24fw, a main-sequence star monitored by the All-Sky Automated Survey for SuperNovae, faded rapidly in late 2024, with the event lasting until June 2025. The pre-dimming SED of ASASSN-24fw indicates that it is a F-type main sequence star with persistent infrared excess having a fractional luminosity of around 0.5%. We modeled this infrared excess using two-component blackbody models and find that the warmer component has a effective temperature of 1072K while the colder component has an effective temperature of 387K. The publicly available survey lightcurves show that ASASSN-24fw is a stable star in brightness and hence the dimming is due to external companion. While long period stellar dimming events have been increasingly discovered over the past decade, what distinguishes the 275 days long event ASASSN-24fw is its unique "flat-base" in the lightcurve, which has a duration of nearly 200 days. In order to discern the physical properties of the occulter and the host star, we have analyzed this event using available photometric and spectroscopic data taken during dimming. In particular, we have fitted two different lightcurve models and performed spectroscopic analysis. Our first modeling method quantifies the parameters of the lightcurve and reveals the multiple phases of ingress, thereby indicating variation in the density of the occulting material. Our second lightcurve model suggests that a gas-giant (perferably a brown dwarf) with a minimum mass of 3.42MJ and having massive (0.17 au) circum-planetary rings. The near-infrared spectra taken during dimming show elevated levels of infrared excess and also indicate that the central object is likely a brown dwarf with a spectral-type of M8. We notice that the H line is transient in nature and has relation with the composition of the occulter. We observed this event using LCOGT to measure the extinction during dimming. We find that the the occulting material may contain large bodies and gas. We also find a companion in the LCOGT images within 3" radius. Our analysis suggests that this is a cool main-sequence star at a separation of 3000 au and unlikely to be un-associated with ASASSN-24fw. Finally, we suggest that systems like ASASSN-24fw are very rare and unique and more follow-up observations are needed to constrain the stellar properties, the circum-stellar environment and the evolutionary mechanisms for such systems.
@@
binary – stellar dimming – circum-stellar disc – exoplanet1 Introduction
@@
Nearly all stars are born surrounded by dense circum-stellar envelopes (Ribas et al., 2014; Manara et al., 2023). These envelopes process the natal proto-stellar cloud into planetary discs over a period of a few to 10s of million years (Williams & Cieza, 2011). These circum-stellar young protoplanetary discs are the birthplace of planetary systems including planetesimal belts (Wyatt et al., 2015; Lovell et al., 2021). Circum-stellar discs manifest in various forms (Hughes et al., 2018) and now we know that they are ubiquitous throughout the pre- to post-main sequence lifetimes of stars (Jewitt et al., 2009; Dennihy et al., 2020; Zhao et al., 2020; Raymond & Nesvorný, 2022; Myers et al., 2024; Swan et al., 2024; Farihi et al., 2025). The dynamical excitation induced erosive or destructive collisions of the planetesimals (e.g. a phase of Heavy-Bombardment period in Solar System (Gomes et al., 2005; Bottke & Norman, 2017)) is a common mechanism to generate detectable levels of dust manifesting as debris discs around stars (e.g. Marshall et al., 2021; Matrà et al., 2025).
@@
Circum-stellar discs significantly change their structure, dynamics, and composition over time and, thus their architectures and brightness are correlated with the stellar age (Cao et al., 2023; Matrà et al., 2025), e.g., a pre-main sequence is surrounded by a protoplanetary disc of mainly gas and dust, the transition or intermediate disc phase has altered composition and geometry while the debris disc is further evolved like the star hosting it and contains mainly of collisional rocky bodies and dust generated by destructive processes (Hughes et al., 2018; Su et al., 2019; Moór et al., 2024). Thus, the detection of such debris discs points towards successful planet-formation (or even destruction) processes.
@@
These circum-stellar discs can be detected through the presence of excess flux from infrared to millimeter wavelengths (Su et al., 2019). As the material in these discs re-emits the incident stellar radiation into longer wavelengths, the composition of the debris disc, structure and its evolutionary status can be inferred from the location of the peak of the excess flux (Wyatt et al., 2015; Hughes et al., 2018; Cao et al., 2023). This information in parallel with the nature of the optical variability can be used to understand the nature of the circum-stellar material e.g., either proto-planetary, transitional, or debris discs. For example, V488 Per has the most infrared excess of any circum-stellar disc and is therefore thought to be surrounded by an Extreme Debris Disc (Moór et al., 2024) which has a fractional luminosity 1%, HD 166191 is surrounded possibly by a transition disc (Kennedy et al., 2014) (although this is still controversial, Garcia & Hughes, 2019), and RZ Piscium contains a debris disc perturbed by a red-dwarf companion (Kennedy et al., 2020). However, there are examples of optical variability without infrared excess e.g. Boyajian’s star (KIC 8452642; Boyajian et al., 2016) had no transient infrared excess (Marengo et al., 2015) and therefore it was attributed to the disintegration of a family of exo-comets (Bodman & Quillen, 2016). Another such event around ASASSN-21qj has likewise been attribted to exocometary activity and exhibited a transient infrared excess (Marshall et al., 2023).
@@
Sometimes, the variability in stellar brightness is also caused by systems containing circum-planetary, or circum-stellar discs. For example, the Mamajek’s object (Mamajek et al., 2012), RW Aur (Rodriguez et al., 2013; Rodriguez et al., 2016), V409 Tau and AA Tau (Rodriguez et al., 2015), "The Random Transitor" type objects HD 139139 (Rappaport et al., 2019b), and EPIC 204376071 (Rappaport et al., 2019a), Gaia21bcv (Hodapp et al., 2024), ASASSN-21js (Pramono et al., 2024). Another example is VVV-WIT-08/10 (Smith et al., 2021), discovered by the VISTA Variables in the Vía-Láctea (VVV; Minniti et al. (2010)) survey, where an eclipse lasting 200 days is attributed to an elliptical disc-like occulter, similar to the proposed ring system of J1407b (Kenworthy & Mamajek, 2015).
@@
One of the important distinguishing points between variability due to circum-stellar material and circum-planetary material is the depth and duration of the dimming. While the variability due to the former is irregular and a combination of several short-period dips, the variability due to the latter could manifest as a dip both larger in depth and period. It may also show some extra features depending on the orientation of the system. As it is not exactly clear how such objects with massive rings form and if there is any correlation between the circum-stellar environment and the properties of such objects, we can build a clearer picture about the such systems only by continuing to identify and analyze such events.
@@
This paper is another attempt to explain the recent stellar dimming event: ASASSN-24fw. This dimming event was reported by ASAS-SN’s transient detection page followed by an Astronomer’s Telegram (ATel No. 16833, JoHantgen et al., 2024). ASASSN-24fw has been the subject of study in two recently published articles by Forés-Toribio et al. (2025) and Zakamska et al. (2025). Our preliminary investigation found that ASASSN-24fw is located close to a star forming region in the Monoceros constellation (Hughes & Baines, 1985; Costado & Alfaro, 2018). However, later we found that both are not related to each other. The star has a persistent infrared excess with a fractional luminosity of 0.5%. We find that a two component blackbody model with effective temperatures of 1072K and 387K can explain the infrared excess in the SED. Since it had a stable optical and mid-infrared pre-dimming brightness, it can be assumed that the circum-stellar material is distributed around the star. In September 2024, it experienced a steep reduction in brightness which lasted until around the first week of October 2024. The star remained around the same low-brightness until the mid-May of 2025 and the egress occurred close to the first week on June 2025. Another Telegram by Nair & Denisenko (2024) (ATel No. 16919) noted similar dimming episodes of this star around the years 1937 and 1982, suggesting a possible periodicity of 16,000 days (or 43.8 years). The drastic drop in the brightness of ASASSN-24fw from g’ 13.20 mag to 16.70 mag (close to the limit of ASAS-SN sensitivity at 17.50 mag) rapidly within a fortnight followed by a nearly stable brightness for six months and mirrored egress is the unique feature of this event.
From the shape and analysis of the lightcurve, we suggest that the dimming actually started around March 2024 and the ingress happened in three stages. We suggest that the dimming was caused by a central body with massive circum-planetary rings ( 0.17 au in size). The central body is a gas-giant with a mass of at least 3.42MJ and the material density in the rings is inversely proportional to the radial distance from it. We think that the gas-giant is likely to be a brown-dwarf although we do not rule out a super-Jupiter interpretation. We present previously unpublished information tracing the evolution of this event after its discovery, seeking to characterize both the size and mass of the star and the occulting body, and identify its size, chemical composition, and mass of the constituent material around it. We also discuss the studies by Forés-Toribio et al. (2025) and (Zakamska et al., 2025) later in our paper from a broader perspective.
@@ 2.1 All-Sky Automated Survey for SuperNovae (ASAS-SN)
The All-Sky Automated Survey for SuperNovae (ASAS-SN) (Shappee et al., 2014) is a long term survey project to monitor the sky down to a limiting magnitude of V 17 mag using a global network of telescopes. The focus of the survey is to find nearby supernovae (SNe) and other transient sources. Currently, this network consists of six fully robotic units on Mount Haleakala in Hawaii, Cerro Tololo in Chile, LCO sites in South Africa and USA, and another in China. Each unit consists of several robotic 14 cm aperture telescopes and is hosted by the corresponding station and each telescope is mounted with a Nikon telephoto lens and a 2k x 2k thinned CCD, giving a 4o.5 x 4o.5 field-of-view and a 7.8" pixel scale. These 24 telescopes together survey entire night sky every night and discoveries are announced within few hours of the data being taken. The data can be accessed through https://asas-sn.osu.edu/.
@@ 2.2 LCOGT
We monitored ASASSN-24fw with time-series photometry from February 22nd to April 5th 2025 (PID:DDT2025A-003, PI: Sarang Shah) using the Las Cumbres Observatory’s network of 1-m telescopes with the SINISTRO imaging instrument (Brown et al., 2013). The stellar brightness was recorded in four filter bands: SDSS g’, r’, i’, and PANSTARRS zs. Integration times were 12s in g’, 18s in r’, 26s in i’, and 94s in zs. This provided a SNR100 in each filter for an occulted magnitude of 16.5. The photometric measurements were taken from the pipeline-produced BANZAI catalog (McCully et al., 2018). If the catalog (or target) was missing from the observatory provided data, the stellar magnitude was determined using aperture photometry performed using the Python-based Photutils package (Bradley et al., 2024) using a 5" radius aperture and a background annulus between 10" and 15".
In the LCOGT images, we find a faint companion star which is also present in Gaia DR3 with > 19 mag and at a separation of 3". In Figure (1), we show a cropped 3"x3" LCOGT g-band image where the colorbar represents the pixel count. ASASSN-24fw and its faint companion are located in the center of the red-circle (and image). If bound to ASASSN-24fw, at a distance of 1kpc, this companion is located at a distance of 3000 au. To check how likely it is to have a chance alignment for a faint companion at a separation 3", we stacked our images to increase the SNR of the companion. We then performed the psf photometry of the companion by masking ASASN24fw. The (g’-r’), (r’-i’), and (i’-z’) colors that we get are 1.0, 0.66, and -0.83 respectively. The former two colors indicate a K-type main-sequence star. The negative (i-z) color is due to bad sky in the z-band images. Further, we find there are 439 sources fainter than or equal to the brightness of ASASSN-24fw. We define the chance alignment probability (P(X)) of a star within 3" radius of ASASN-24fw as:
where the FoV is 13.2 x 13.2 sq. arcmin and is the area of a circular region of 3" radius around ASASN-24fw. We find that P(X) is only 2% which indicates that there is a high chance that the companion might not be a chance alignment. But this can be only confirmed in future by conducting follow-up campaign to study the proper motion of this faint companion.
@@ 2.3 AAVSO
American Association of Variable Star Observers (AAVSO) (Kloppenborg, 2025) is a association of professional and amateur astronomers that observe variability related phenomena and catalog their high quality photometric observations. This association of professional and amateur astronomers regularly observes and analyses variable stars and their observers are located across the globe. We obtained the AAVSO observed data from their website111https://www.aavso.org/data-download/.
@@ 2.4 NEOWISE
The NEOWISE mission (WISE Team, 2020) utilized the Wide-field Infrared Survey Explorer spacecraft and a survey strategy similar to WISE mission (Wright et al., 2010) to scan the sky between the ecliptic poles in great semi-circles and nearly perpendicular to the Earth- Sun line. This survey strategy created a set of multiple independent exposures on each point of the sky over the 10 year period and had observed the region including ASASSN-24fw in the past decade. Thus, we could obtain and combine all the single exposure NEOWISE photometry which is an array of one observation every six months. We found no infrared variability of ASASSN-24fw in the NEOWISE lightcurve as all NEOWISE observations were taken outside the eclipse.
@@ 2.6 Spectral Observations
We obtained a low resolution spectrum with Keck LRIS instrument (Oke et al., 1995; Rockosi et al., 2010) on 29 October 2024. The resolution of the grating element was 1300 and the exposure time was 2700s. We also obtained a second high resolution transmission spectrum using the HiRES instrument on 28 January 2025. The resolution of the grating element that was used was 25,000 and the exposure time was 600s. Both of these observations were conducted when the star was dimmed and were extracted using Pypelt (prochaska2020). The wavelengths were then calibrated using a combination of Th-Ar spectra and telluric absorption with the standard star observations using xtellcor (Vacca et al., 2003). We also obtained a NIR low resolution spectrum (R 3000) using the TripleSpec instrument (Herter et al., 2008) at the 5.1m Hale Telescope at Palomar Observatory on 21 December 2024. The spectrum was extracted using the IDL package spextool (Cushing et al., 2004). The extracted spectra were also flux calibrated and corrected for telluric absorption in a similar way.

### v2

@@ Abstract
ASASSN-24fw, monitored by the All-Sky Automated Survey for SuperNovae, underwent a rapid and deep dimming event beginning in late 2024 and lasting until June 2025. The pre-dimming spectral energy distribution indicates that ASASSN-24fw is an F-type main-sequence star with a persistent infrared excess, corresponding to a fractional luminosity of about 10 per cent. Public survey lightcurves show that the star is otherwise photometrically stable, implying that the dimming was caused by an external occulter. While long-duration stellar dimming events have become increasingly common in recent years, ASASSN-24fw is distinguished by a pronounced flat-bottomed lightcurve lasting nearly 200 days. We analyse this event using available photometric and spectroscopic data obtained during the dimming phase. Two independent lightcurve models are applied. The first characterizes the multiple phases of the ingress structure, while the second model suggests that the occulter is a gas-giant companion, likely a brown dwarf, with a minimum mass of 3.42 MJ and an extended ring system with a radial extent of approximately 0.17 au. Near-infrared spectra obtained during dimming show enhanced infrared excess and are consistent with an M8-type brown dwarf companion. Optical spectra reveal variable H emission during the event. Post-dimming spectra show no evidence for ongoing accretion, suggesting that the H variability is associated with the occulter rather than the host star. We conclude that ASASSN-24fw represents a rare and unusual system, and that further observations are required to better constrain its stellar properties, circumstellar environment, and evolutionary history.
@@
binary – stellar dimming – circumstellar disc – exoplanet1 Introduction
@@
Nearly all stars are born surrounded by dense circumstellar envelopes (Ribas et al., 2014; Manara et al., 2023). These envelopes process the natal proto-stellar cloud into planetary discs over a period of a few to 10s of million years (Williams and Cieza, 2011). These circumstellar young protoplanetary discs are the birthplace of planetary systems including planetesimal belts (Wyatt et al., 2015; Lovell et al., 2021). circumstellar discs manifest in various forms (Hughes et al., 2018) and now we know that they are ubiquitous throughout the pre- to post-main sequence lifetimes of stars (Jewitt et al., 2009; Dennihy et al., 2020; Zhao et al., 2020; Raymond and Nesvorný, 2022; Myers et al., 2024; Swan et al., 2024; Farihi et al., 2025). The dynamical excitation induced erosive or destructive collisions of the planetesimals (e.g. the proposed Late Heavy Bombardment period in the Solar System Gomes et al. (2005); Bottke and Norman (2017)) is a common mechanism to generate detectable levels of dust manifesting as debris discs around stars (e.g. Marshall et al., 2021; Matrà et al., 2025).
@@
Circumstellar discs significantly change their structure, dynamics, and composition over time and, thus their architectures and brightness are correlated with the stellar age (Cao et al., 2023; Matrà et al., 2025), e.g., a pre-main sequence is surrounded by a protoplanetary disc of mainly gas and dust, the transition or intermediate disc phase has altered composition and geometry while the debris disc is further evolved like the star hosting it and contains mainly of collisional rocky bodies and dust generated by destructive processes (Hughes et al., 2018; Su et al., 2019; Moór et al., 2024). Thus, the detection of such debris discs points towards successful planet-formation (or even destruction) processes.
@@
These circumstellar discs can be detected through the presence of excess flux from infrared to millimeter wavelengths (Su et al., 2019). As the material in these discs re-emits the incident stellar radiation into longer wavelengths, the composition of the debris disc, structure and its evolutionary status can be inferred from the location of the peak of the excess flux (Wyatt et al., 2015; Hughes et al., 2018; Cao et al., 2023). This information in parallel with the nature of the optical variability can be used to understand the nature of the circumstellar material e.g., either proto-planetary, transitional, or debris discs. For example, V488 Per has the most infrared excess of any circumstellar disc and is therefore thought to be surrounded by an Extreme Debris Disc (Moór et al., 2024) which has a fractional luminosity 1%, HD 166191 is surrounded possibly by a transition disc (Kennedy et al., 2014) although this is still controversial (see e.g. Garcia and Hughes, 2019), and RZ Piscium contains a debris disc perturbed by a red-dwarf companion (Kennedy et al., 2020). However, there are examples of optical variability without infrared excess e.g. Boyajian’s star (KIC 8452642; Boyajian et al., 2016) had no transient infrared excess (Marengo et al., 2015) and therefore it was attributed to the disintegration of a family of exo-comets (Bodman and Quillen, 2016). Another such event around ASASSN-21qj has likewise been attribted to exocometary activity and exhibited a transient infrared excess (Marshall et al., 2023).
@@
Sometimes, the variability in stellar brightness is also caused by systems containing circumplanetary, or circumstellar discs. For example, Mamajek’s object (Mamajek et al., 2012), RW Aur (Rodriguez et al., 2013, 2016), V409 Tau and AA Tau (Rodriguez et al., 2015), "The Random Transitor" type objects HD 139139 (Rappaport et al., 2019a), and EPIC 204376071 (Rappaport et al., 2019b), Gaia21bcv (Hodapp et al., 2024), ASASSN-21js (Pramono et al., 2024). Another example is VVV-WIT-08/10 (Smith et al., 2021), discovered by the VISTA Variables in the Vía-Láctea (VVV; Minniti et al., 2010) survey, where an eclipse lasting 200 days is attributed to an elliptical disc-like occulter, similar to the proposed ring system of J1407b (Kenworthy and Mamajek, 2015).
@@
One of the important distinguishing points between variability due to circumstellar material and circumplanetary material is the depth and duration of the dimming. While the variability due to the former is irregular and a combination of several short-period dips, the variability due to the latter could manifest as a dip both larger in depth and period. It may also show some extra features depending on the orientation of the system. As it is not exactly clear how such objects with optically thick rings form and if there is any correlation between the circumstellar environment and the properties of such objects, we can build a clearer picture about the such systems only by continuing to identify and analyze such events.
@@
This paper is another attempt to explain the recent stellar dimming event: ASASSN-24fw. This dimming event was reported by ASAS-SN’s transient detection page followed by an Astronomes Telegram (ATel No. 16833, JoHantgen et al., 2024). ASASSN-24fw has been the subject of study in two recently published articles by Forés-Toribio et al. (2025) and Zakamska et al. (2025). In September 2024, it experienced a steep reduction in brightness which lasted until around the first week of October 2024. The star remained around the same low-brightness until the mid-May of 2025 and the egress occurred close to the first week on June 2025. Another Telegram by Nair and Denisenko (2024) (ATel No. 16919) noted similar dimming episodes of this star around the years 1937 and 1982, suggesting a possible periodicity of 16,000 days (or 43.8 years). The drastic drop in the brightness of ASASSN-24fw from 13.20 mag to 16.70 mag (close to the limit of ASAS-SN sensitivity at 17.50 mag) rapidly within a fortnight followed by a nearly stable brightness for six months and mirrored egress is the unique feature of this event.
@@ 2.1 All-Sky Automated Survey for SuperNovae (ASAS-SN)
The All-Sky Automated Survey for SuperNovae (ASAS-SN) (Shappee et al., 2014) is a long term survey project to monitor the sky down to a limiting magnitude of V 17 mag using a global network of telescopes. The focus of the survey is to find nearby supernovae (SNe) and other transient sources. Currently, this network consists of six fully robotic units on Mount Haleakala in Hawaii, Cerro Tololo in Chile, LCO sites in South Africa and USA, and another in China. Each unit consists of several robotic 14 cm aperture telescopes and is hosted by the corresponding station and each telescope is mounted with a Nikon telephoto lens and a 2k x 2k thinned CCD, giving a 4545 field-of-view and a 78 pixel scale. These 24 telescopes together survey entire night sky every night and discoveries are announced within few hours of the data being taken. The data can be accessed through https://asas-sn.osu.edu/.
@@ 2.2 LCOGT
We monitored ASASSN-24fw with time-series photometry from February 22nd to April 5th 2025 (PID:DDT2025A-003, PI: Sarang Shah) using the Las Cumbres Observatory’s network of 1-m telescopes with the SINISTRO imaging instrument (Brown et al., 2013). The stellar brightness was recorded in four filter bands: SDSS , , , and PANSTARRS zs. Integration times were 12s in , 18s in , 26s in , and 94s in zs. This provided a SNR100 in each filter for an occulted magnitude of 16.5. The photometric measurements were taken from the pipeline-produced BANZAI catalog (McCully et al., 2018). If the catalog (or target) was missing from the observatory provided data, the stellar magnitude was determined using aperture photometry performed using the Python-based Photutils package (Bradley et al., 2024) using a 5 radius aperture and a background annulus between 10 and 15.
In the LCOGT images, we find a faint companion star which is also present in Gaia DR3 with > 19 mag and at a separation of 3. In Figure (1), we show a cropped 33 LCOGT -band image where the colorbar represents the pixel count. ASASSN-24fw and its faint companion are located in the center of the red-circle (and image). If bound to ASASSN-24fw, at a distance of 1kpc, this companion is located at a distance of 3000 au. To check how likely it is to have a chance alignment for a faint companion at a separation 3, we stacked our images to increase the SNR of the companion. We then performed the psf photometry of the companion by masking ASASSN-24fw. The (-), (-), and (-) colors that we get are 1.0, 0.66, and -0.83 respectively. The former two colors indicate a K-type main-sequence star. The negative (-) color is due to bad sky in the -band images. Further, we find there are 439 sources fainter than or equal to the brightness of ASASSN-24fw. We define the chance alignment probability () of a star within 3 radius of ASASSN-24fw as:
where the FoV is 13.2 13.2 sq. arcmin and is the area of a circular region of 3 radius around ASASSN-24fw. We find that () is only 2% which indicates that there is a high chance that the companion might not be a chance alignment. But this can be only confirmed in future by conducting follow-up campaign to study the proper motion of this faint companion.
@@ 2.3 AAVSO
American Association of Variable Star Observers (AAVSO) (Kloppenborg, 2025) is an association of professional and amateur astronomers that observe variability related phenomena and catalog their high quality photometric observations. This association of professional and amateur astronomers regularly observes and analyses variable stars and their observers are located across the globe. We obtained the AAVSO observed data from their website111https://www.aavso.org/data-download/.
@@ 2.4 NEOWISE
The NEOWISE mission (WISE Team, 2020) utilized the Wide-field Infrared Survey Explorer (WISE) spacecraft and a survey strategy similar to WISE mission (Wright et al., 2010) to scan the sky between the ecliptic poles in great semi-circles and nearly perpendicular to the Earth-Sun line. This survey strategy created a set of multiple independent exposures on each point of the sky over the 10 year period and had observed the region including ASASSN-24fw in the past decade. Thus, we could obtain and combine all the single exposure NEOWISE photometry which is an array of one observation every six months. We found no infrared variability of ASASSN-24fw in the NEOWISE lightcurve as all NEOWISE observations were taken outside the eclipse.
@@ 2.6 Spectral Observations
We obtained a low resolution spectrum with Keck LRIS instrument (Oke et al., 1995; Rockosi et al., 2010) on 29 October 2024. The resolution of the grating element was 1300 and the exposure time was 600s. We also obtained a second high resolution transmission spectrum using the HiRES instrument on 28 January 2025. The resolution of the grating element that was used was 25,000 and the exposure time was 2700s. Both of these observations were conducted when the star was dimmed and were extracted using Pypelt (Prochaska et al., 2020). The wavelengths were then calibrated using a combination of Th-Ar spectra and telluric absorption with the standard star observations using xtellcor (Vacca et al., 2003). We also obtained a single NIR low resolution spectrum (R 3000) using the TripleSpec instrument (Herter et al., 2008) at the 5.1m Hale Telescope at Palomar Observatory on 21 December 2024. The exposure time was 120s and the spectrum was extracted using the IDL package spextool (Cushing et al., 2004). The extracted spectrum was also flux calibrated and corrected for telluric absorption in a similar way. We recently obtained two post-dimming optical spectra of ASASSN-24fw on 13 October 2025 using the MIKE spectrograph on Magellan Telescope. The exposure time of each spectrum was 600s with a effective resolution of 57,000 and they were reduced using Daniel Kelson’s pipeline available at the Carnegie Observatories Software Repository222http://code.obs.carnegiescience.edu/. This was taken to compare the during-dimming and out-of dimming nature of the star. We show this raw stacked spectrum in Figure (5) of Appendix (A).